OpenStudy (toxicsugar22):

how do you solve for r: A=P(1+r)^2

OpenStudy (anonymous):

divide by P, take the square root, subtract 1. i assume r is positive

OpenStudy (anonymous):

a=p(1+r)^2 a/p=(1+r)^2 => r=1+(a/p)^2 (or) r=1-(a/p)^2 :)

OpenStudy (anonymous):

bracket it is square root, not square

OpenStudy (anonymous):

\[A=P(1+r)^2\implies(1+r)^2=\frac{A}{P}\implies1+r=\sqrt{\frac{A}{P}}\implies r= \sqrt{\frac{A}{P}}-1\]

OpenStudy (anonymous):

what bendt said!

myininaya (myininaya):

very pretty bendt

OpenStudy (anonymous):

Thanks :P

OpenStudy (anonymous):

although in reality it should be \[\pm\sqrt{\frac{A}{P}}\]

OpenStudy (toxicsugar22):

ok thanks

OpenStudy (anonymous):

but in the application of this formula it is really just the positive root that counts.

myininaya (myininaya):

that is right bu what does A stand for?

OpenStudy (anonymous):

A stands for the thing on the right!

myininaya (myininaya):

i know that lol but it looks like thats a formula for that business crap stuff

OpenStudy (anonymous):

An area?

OpenStudy (anonymous):

yea rite sattelite73

OpenStudy (anonymous):

i have to go but i am not leaving until myininaya compliments my icon

myininaya (myininaya):

like r is the rate and P is the principal

OpenStudy (anonymous):

A is amount if you invest P dollars at r rate of interest for 2 years

myininaya (myininaya):

i dont know how i feel about your icon yet so you won't leave until i have a feeling i guess so you can lose money? so A can be negative so you can have + or - on that one thing

OpenStudy (anonymous):

the negative root makes no sense in this i don't think. ok now i am going to change my icon since i did not get your seal of approval.

myininaya (myininaya):

lol

myininaya (myininaya):

you should blow me away satellite

myininaya (myininaya):

with your new icon

myininaya (myininaya):

you can't go wrong with icon that says myininaya is the best!

OpenStudy (anonymous):

satellite is that icon one of mondrian's paintings?